OpenStudy (anonymous):

Write each function in vertex form. Y=2x^2+x Y=-2x^2-4x+6 Y=x^2-2x-6

OpenStudy (zpupster):

there are a couple of methods but i use this it seems easier but you have to memorize -b/2a this is from the form ax^2 + bx + c so from the first equation b=1 and a=2 so - 1/2(2) = -1/4 now we have x(x,?) insert x back into the eqUATION y=2(-1/4)^2+(-1/4) y=2(1/16)-1/4=1/8-2/8=-1/8 so (-1/4,-1/8) is our x and y or our h and k because the vertex form is: y = a(x - h)² + k so plug in our values y=2(x+1/4)^2 + (-1/8) or y=2(x+.25)^2 - .125 see graph

OpenStudy (zpupster):

i have time i can show you another method on the second eq. and then you can pick a method and do the third. this method we will complete the square first lets factor out coefficient on a to get ax^2 + bx + c Y=-2x^2-4x +6 coefficient is -2 so y=-2(x^2 + 2x + ? ) +6 ?=1 putting 1 here completes our square y= -2((x+1)(x+1)) +6 so we have y= -2(x+1)^2 ?? +6 notice that we added 1 to complete the square here it gets tricky because to keep everything even we have to subtract 1 and it is not really 1 we have to remember the -2 coefficient so -2(1) is -2 we have to add 2 to keep the equation even. so y= -2(x+1)^2 +2 +6 y= -2(x+1)^2 +8 y = a(x - h)² + k our (h,k) is (-1,8) check with the graph and ok you do 3) hope this helps!!!